How do you solve for the exact values if you're given a problem like, sec[arcsin(-5/13)]
Answer by stanbon(75887)

How do you solve for the exact values if you're given a problem like,
sec[arcsin(-5/13)]
------
You are told the sine of an angle is -5/13
That angle must be in QIII or in QIV where y is negative.
------
Since the sine = y/r = -5/13, y = -5 and r = 13
Then x = sqrt[13^2-5^2] = +12 or -12
--------
Then, since secant = r/x, sec[arcsin(-5/13)] = 13/12 in QIV
or sec[arcsin(-5/13)] = -13/12 in QIII
------
Cheers,
Stan H.